# Perturbation-based Effect Measures for Compositional Data

### v3

@@ Acknowledgements
We first note that
$\widetilde{\tau}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
since $\widehat{p}\leq 1$. Similarly, we have that
Therefore, by several applications of Lemma S2, we have that
S3.3 Estimating differentiable regression functions
$p(\ell)=\beta^{1}+\beta^{2}(\ell-L_{i})+\beta^{3}(\ell-L_{i})^{2}$
3:Extract weight matrix $K$ from $\widehat{\mu}$.
4:for $i=1,\ldots,n$ do
6:     Set $\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)$.
7:     Set $\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i)$.
S3.4 Location-scale modelling of the score function
$L=m(W)+v(W)^{1/2}\zeta,$
Algorithm S3 Location-scale score modelling estimation
3:Regress $L$ on $W$ using $I_{1}$ yielding regression estimate $\widehat{m}$.
7:return $\widehat{\rho}$.
Section S4.1 provides the details behind the experiments in Table 1 in Section 1. Section S4.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional. Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.1.1. Section S4.4 includes an additional figure showing the semi-synthetic dataset used in Section 4.1.1. Section S4.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.1.2.
S4.1 Simulations from Section 1
In this section we describe the data-generating processes and methods used in the numerical illustrations of Example 1 and Example 2. The results of are provided in Table 1.
S4.1.1 Example 1: Effects of absence or presence of microbes
$Z:=(Z^{1},W^{2}(1-Z^{1}),(1-W^{2})(1-Z^{1}))\in\Delta^{2},$
$\mathrm{CKE}^{1}$ is computed using the $\lambda$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
S4.1.2 Example 2: Effects of increased diversity
$\mathrm{CDI}_{\mathrm{Gini}}$ is computed using the $\tau$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
S4.2 Semiparametric estimators
  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (14) and (15). See Algorithms 1 and 2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
  • NPM_no_x: The same estimators as NPM but without cross-fitting.
  • PLM: The partially least model based estimators in (17) including cross-fitting. See Algorithm 3 with $K=2$ for the detailed estimators.
  • PLM_no_x: The same estimators as PLM but without cross-fitting.
  • Binary $L$ with partially linear $Y$:
  • Binary $L$ with nonparametric $Y$:
  • Continuous $L$ with partially linear $Y$:
    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
  • Continuous $L$ with nonparametric $Y$:
S4.3 Description of the construction of groups in Section 4.1.1
$(\widehat{\pi}_{E}(0),\widehat{\pi}_{E}(1))$ $=(0.454,0.546),$
$(\widehat{\pi}_{C}(0),\widehat{\pi}_{C}(1))$ $=(0.761,0.239),$
$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{E}^{1}(\texttt{education}_{i}),$
and
$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{E}^{2}(\texttt{education}_{i}),$
and
where $p_{R}^{1}$ is as in (S8) and $p_{C}^{1}$ is a probability vector satisfying
where $p_{R}^{2}$ is as in (S9) and $p_{C}^{2}$ is a probability vector satisfying
S4.4 Additional figure for the ‘Adult’ data experiment in Section 4.1.1
Figure S7 complements Figure 4 in Section S4.4 by summarizing the semi-synthetic ‘Adult’ dataset when grouping on compensation.
Figure S7: Summary of the semi-synthetic ‘Adult’ dataset when grouping on compensation. Left: Racial compositions in a ternary plot. Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red). Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle). Depending on the color the dependence between $Y$ and the Gini coefficient changes. Note that the Gini coefficient represents a similarity rather than a diversity measure.
S4.5 Semi-synthetic individual-level experiment using groups from Section 4.1.1
In this section we consider a variant of the experiment performed in Section 4.1.1 where instead of aggregating the full dataset into groups, we preserve the individual observations but generate new responses based on the groups. More specifically, we generate the groups as before and estimate the probability that $Y=1$ within each group and use this probability to generate new $Y$s for each observation. For each individual, we thus have a response $Y$, a race $R$, a ‘contextual’ variable $Z$ that gives the racial distribution of each group and additional covariates $X$. By construction of the groups and sampling of the $Y$, we know that there is a positive effect of diversity on the income. We now apply the naive_diversity and $\mathrm{CDI}_{\mathrm{Gini}}$ methods as in Section 4.1.1 except we always control for the race of inviduals and omit the unstable estimators that condition on $Z$. The results can be seen in Table 5.
Method Grouping on compensation
Estimate $95\%$ CI
[…]
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.487}}$ $(0.282,0.692)$
Table 5: Table of estimates in the individual-level ‘Adult’ data experiment. Bold red numbers indicate a significant negative effect of diversity at a 5% level, while bold green indicates a significant positive effect. The naive diversity method indicates that there is a significant negative effect of diversity on compensation, even when controlling for individual race ($R$) and additional covariates ($X$), despite the fact that the data is constructed with a positive effect of diversity. The CDI correctly identifies that the effect is positive.
S4.6 Description of additional control variables used in Section 4.1.2
School income estimate
School grade range (grade referring to years of school)
Student attendance rate
% chronically absent students
Economic Need Index
Community school
An indicator of whether the school is a community school.
% English language learners
Percentage of students who are learning English for the first time.
Rigorous Instruction %
Collaborative Teachers %
Supportive Environment %
Effective School Leadership %
Strong Family-Community Ties %
Trust %
S4.7 Additional plots from gut microbiome analysis in Section 4.2
In this section we provide an additional figure showing our results from the analysis of the gut microbiome data in Section 4.2. In Figure S8, we plot the most significant hits for $\mathrm{CFI}{\mathrm{unit}}$ and $\mathrm{CKE}$ both estimated using random forests together with $\mathrm{CFI}_{\mathrm{mult}}$ estimates. We plot the estimates scaled by the estimated standard deviation of $L$ to put the effects on the same scale (we do not correct the confidence intervals for the estimation of this standard deviation). Furthermore, we flip the sign of $\mathrm{CKE}$ to keep the interpretation similar to $\mathrm{CFI}$, that is, that we are looking at the effect of increasing $Z^{j}$.
Appendix S5 Proof of Theorem 1
S5.1 Proof of (a)
@@
By the Cauchy–Schwarz inequality, Assumption 1 (f) and Lemma S3, we have
@@
Combining the above with Assumption 1 (a) and (e), we have
@@
By the above, Assumption 1 (c), (d) and (a), Lemma S2 (a) and Jensen’s inequality, we have that
@@
Assumption 1 (c), (d) and (a) and Lemma S3 thus yield that
@@
Therefore, using Assumption 1 (a), the Cauchy–Schwarz inequality and Jensen’s inequality, we have
@@
The conclusion now follows from a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20] and Lemma S4 (a), which we can apply by Assumption 1 (a) and (b), since
@@
Applying the triangle and Cauchy–Schwarz inequality, followed by Assumption 2 (b) and (e) leads to
@@
Combining Assumption 2 (b), (c) and (a), we conclude that
@@
Therefore, using Assumption 2 (a)
@@
Using this, Assumption 2 (b), (S15) and (S16), we have by Lemma S2 that
@@
By Assumption 3 (c), we have $(\textup{I})=o_{\mathcal{P}}(N_{k}^{-1/2})$. Since the summands of the third term are conditionally i.i.d. and mean zero given $\widehat{m}_{k}$ and $(W_{i})_{i\in I_{k}}$, we have by the conditional version of Chebyshev’s inequality that
@@
Combining the above with Assumption 3 (a) and (c), we conclude that
@@
by Assumption 3 (a) and (b). Thus, for all $k\in[K]$, we note that, using the fact that $Y-g_{P}(W)=\varepsilon+\theta_{P}\xi$ and (S21),
@@
By the Cauchy–Schwarz inequality and Assumption 3 (f), we have
@@
Next, combining (S21), Chebyshev’s inequality and Assumption 3 (a) yields that
@@
hence Assumption 3 (b) and Lemma S2 (d) yields
@@
An analogous argument can be used to show that $\widetilde{(\textup{III})}_{g}=o_{\mathcal{P}}(1)$. Using $\sum_{i}a_{i}^{2}a_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}a_{i}^{2}\right)$ and Assumption 3 (f), we have
@@
so by Lemma S2, (S25) and Assumption 3 (a) and (b), (S26) holds.
Finally, combining (S22), (S26) with Lemmas S3 and S2 using Assumption 3 (a) and (b), we have that
@@
The conclusion now follows from Assumption 3 (a) and (b), Lemma S4 (b) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
S5.4 Auxiliary results
This is a collection of minor or known results that we have included for completeness.
@@
Appendix S6 Miscellaneous results
The following result provides a relationship between the $\mathrm{CFI}^{j}$ and coefficients in a well-specified log-contrast regression model. The result has been previously proven in Huang et al. [2023].
@@
$\mathrm{CFI}^{j}_{\mathrm{mult}}=\beta^{j}.$
@@
We thus have
and therefore
We conclude that
The next result shows that the effects proposed in Cox [1971] are special cases of average perturbation effects in a parametric model.

### v4

@@ Acknowledgements
We have
and therefore
We conclude that
Appendix S3 Perturbations as causal quantities
Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions. To make this distinction clear, we discuss here how our perturbation framework can be used for causal inference. Causal models [Pearl, 2009, Rubin, 2005] provide a mathematical framework for modeling changes to the data-generating mechanism – often called interventions. In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data-generating mechanism under hypothetical interventions. The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution. Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism. To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework [Pearl, 2009, Bongers et al., 2021]. Similar considerations, however, also apply with the potential outcome framework.
As a first example, assume that the SCM given in Figure S2 (left) generates the data. Given a perturbation $\psi$ that we are interested in, a derivative-isolating reparametrization (see Section 2.1.2) $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure S2 (right)).
(a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
$Y$ $\leftarrow f(Z,\eta)$
$Y$$Z$$\psi$$f$
bijection $\phi$$\phi(Z)=(L,W)$$Z=\phi^{-1}(L,W)$
$Y$$L$$W$$\psi$$h$$m$$h$
$L$ $\leftarrow m(W,\xi)$
$Y$ $\leftarrow h(L,W,\varepsilon)$
Figure S2: (a) Data-generating SCM on the $Z$-scale. $Z\in\Delta^{d-1}$ is a compositional vector, $Y$ is a response variable and $\eta$ is exogenous noise. (b) Reparametrized SCM on the $(L,W)$-scale. $(L,W)=\phi(Z)$ are the reparametrized predictors, $Y$ is the same response as in (a) and $\xi$ and $\varepsilon$ are exogenous noise variables. The reparametrization needs to disentangle the perturbation effect on $Z$ in a way such that intervening on $Z$ by setting it to $\psi(Z,\gamma)$ corresponds to intervening on only $L$ (red double head arrows).
As a second example, consider the SCM over the variables $(H,Z,Y)\in\mathbb{R}^{q}\times\Delta^{d-1}\times\mathbb{R}$ given by
Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$. One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation [Pearl, 2009], this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$. For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components. Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions. Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors. More specifically, let $\psi$ be a perturbation as defined in Definition 1. Then, for all $\gamma$ sufficiently small, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$. In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$. Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself. The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
and the causal average directional perturbation effect
$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}\big{[}\mathbb{E}[Y\mid Z=\psi(Z,\gamma),H]\big{]},$ (S6)
and hence is identifiable. The formula in (S6) is a special case of the adjustment formula [e.g., Pearl, 2009], but other ways of identifying this effect also exist (e.g., propensity score matching). It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$. In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$. This is for example done in the numerical experiments in Section 4.1.
There are several possible assumptions which permit us to conclude that the causal and non-causal perturbation effect coincide. Obviously, one sufficient condition is to assume that
$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}[Y\mid Z=\psi(Z,\gamma)].$ (S7)
Such a condition holds by definition whenever $Z$ is an exogenous variable (i.e., does not depend on other variables in its structural assignment) but can also hold more generally. A common assumption to ensure (S7) is to assume that $Z$ and $Y$ are not confounded, that is, $Z$ and $Y$ have no common causes (this is known as ignorability in the potential outcome framework). In the SCM $\mathcal{M}$ this would for example mean that $H$ does not exist or either $g$ or $h$ do not depend on $H$. In many cases, this is not satisfied and one should instead carefully argue for which causal assumptions are reasonable and adjust the estimation accordingly, e.g., by conditioning on additional covariates as in (S6).
Appendix S4 Inference and algorithms for perturbation effects
In this section we provide details on the theory and algorithms for the estimation of perturbation effects. Section S4.1 contains a summary of the theoretical results and algorithms for the main estimation procedures. Section S4.2 contains a proof of our main theoretical result. Section S4.3 investigates the robustness of the different estimators. Section S4.4 describes how to modify the estimation of $\tau_{\psi}$ using a derivative-isolating reparametrization in the presence of observations with $\omega_{\psi}(Z)=0$. Section S4.5 describes our procedure for the estimation of the derivative of a generic regression function. Section S4.6 describes our procedure for the estimation of the conditional score function $\rho$ that is required for the estimation of an average partial effect.
S4.1 Theoretical analysis of estimators
In this section we present a result on asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (12). These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere [e.g., Klyne and Shah, 2023, Chernozhukov et al., 2018]. We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.
We start from a given perturbation $\psi$ and a corresponding reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ which reparametrizes $Z$ into $(L,W):=(\phi^{L}(Z),\phi^{W}(Z))$ and assume that $(L,W)$ satisfy (2) if $\psi$ is a binary perturbation and that $\phi$ is derivative-isolating if $\psi$ is a directional perturbation. Our goal is now to use the estimation procedures discussed in Section 3 to estimate the average perturbation effects (i.e., $\tau_{\psi}$ or $\lambda_{\psi}$) consistently and construct asymptotically valid confidence intervals for the estimates. To construct asymptotically valid confidence intervals, we estimate both the target parameter and its asymptotic variance and then show that the resulting standardized estimators are asymptotically standard normal with a uniform convergence rate over the class of possible distributions of the data. Distributionally uniform results are desirable here as they ensure uniform coverage guarantees of the confidence intervals. The following result shows uniform asymptotic normality of each of the three estimation procedures.
Theorem S1 (Uniform asymptotic normality).
  1. (a)
    Nonparametric directional effect estimation: Assumption S1 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\tau},\widehat{\sigma}^{2}_{\tau})$ from Algorithm S1.
  2. (b)
    Nonparametric binary effect estimation: Assumption S2 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\lambda},\widehat{\sigma}^{2}_{\lambda})$ from Algorithm S2.
  3. (c)
    Effect estimation using a partially linear model: Assumption S3 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\theta},\widehat{\sigma}^{2}_{\theta})$ from Algorithm S3.
The result is proven in Section S4.2 below. In the following three sections, we go over each of the settings (a), (b) and (c) individually and provide the required assumptions and detailed algorithms for the target parameter and variance. For the remainder of this section, we omit $\psi$-subscripts for simplicity and use the following notation to help us express the uniformly asymptotic results rigorously. Let $(X_{P,n})_{n\in\mathbb{N},P\in\mathcal{P}}$ denote a family of sequences of random variables indexed by $\mathcal{P}$. For a nonnegative sequence $(a_{n})_{n\in\mathbb{N}}$ we write $X_{P,n}=o_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ it holds that
S4.1.1 Nonparametric estimation for directional effects
We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem S1 (a). A high-level overview of this estimation was given in Section 3.2.2. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S4.6). We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm S1 using a user-specified number of splits $K$ in the cross-validation.
Algorithm S1 Nonparametric estimation of $\tau$
4:for $k=1,\dots,K$ do
6:     Fit nuisance function estimates $\widehat{\eta}_{k}$ using $I_{-k}$.
7:     for $r=0,1$ do
To obtain the result in Theorem S1 (a), we require that the estimation methods $\widehat{f}$ and $\widehat{\rho}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption S1.
  1. (a)
  2. (b)
  3. (c)
  4. (d)
  5. (e)
  6. (f)
S4.1.2 Nonparametric estimation for binary effects
We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem S1 (b). A high-level overview of this estimation was provided in Section 3.2.1. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$. We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm S2 using a user-specified number of splits $K$ in the cross-validation.
Algorithm S2 Nonparametric estimation of $\lambda$
3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
9:Set $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}1-L_{i}$.
10:Set $\widehat{\kappa}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}$.
To obtain the result in Theorem S1 (b), we require that the estimation methods $\widehat{f}$ and $\widehat{\pi}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption S2.
  1. (a)
  2. (b)
  3. (c)
  4. (d)
  5. (e)
S4.1.3 Partially linear model for binary/directional effects
Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem S1 (c). A high-level overview of this estimation was provided in Section 3.2.3. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively. We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm S3 using a user-specified number of splits $K$ in the cross-validation.
Algorithm S3 Partially linear estimation of $\theta$ (where $\theta=\lambda$ or $\theta=\tau$)
3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
9:Set $\widehat{J}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{J}_{k}$
11:for $k=1,\dots,K$ do
To obtain the result in Theorem S1 (c), we require that the estimation methods $\widehat{m}$ and $\widehat{g}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption S3.
$\mathbb{E}_{P}[Y\,|\,L,W]=\theta_{P}L+h_{P}(W).$
  1. (a)
  2. (b)
  3. (c)
  4. (d)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{4}=o_{\mathcal{P}}(1)$
  5. (e)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{g}(W_{i})-g_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
  6. (f)
S4.2 Proof of Theorem S1
In this section, we prove Theorem S1 and some additional auxiliary results.
S4.2.1 Proof of (a)
@@
By the Cauchy–Schwarz inequality, Assumption S1 (f) and Lemma S3, we have
@@
Combining the above with Assumption S1 (a) and (e), we have
@@
By the above, Assumption S1 (c), (d) and (a), Lemma S2 (a) and Jensen’s inequality, we have that
@@
Assumption S1 (c), (d) and (a) and Lemma S3 thus yield that
@@
Therefore, using Assumption S1 (a), the Cauchy–Schwarz inequality and Jensen’s inequality, we have
@@
The conclusion now follows from a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20] and Lemma S4 (a), which we can apply by Assumption S1 (a) and (b), since
@@
Applying the triangle and Cauchy–Schwarz inequality, followed by Assumption S2 (b) and (e) leads to
@@
Combining Assumption S2 (b), (c) and (a), we conclude that
@@
Therefore, using Assumption S2 (a)
@@
Using this, Assumption S2 (b), (S13) and (S14), we have by Lemma S2 that
@@
By Assumption S3 (c), we have $(\textup{I})=o_{\mathcal{P}}(N_{k}^{-1/2})$. Since the summands of the third term are conditionally i.i.d. and mean zero given $\widehat{m}_{k}$ and $(W_{i})_{i\in I_{k}}$, we have by the conditional version of Chebyshev’s inequality that
@@
Combining the above with Assumption S3 (a) and (c), we conclude that
@@
by Assumption S3 (a) and (b). Thus, for all $k\in[K]$, we note that, using the fact that $Y-g_{P}(W)=\varepsilon+\theta_{P}\xi$ and (S19),
@@
By the Cauchy–Schwarz inequality and Assumption S3 (f), we have
@@
Next, combining (S19), Chebyshev’s inequality and Assumption S3 (a) yields that
@@
hence Assumption S3 (b) and Lemma S2 (d) yields
@@
An analogous argument can be used to show that $\widetilde{(\textup{III})}_{g}=o_{\mathcal{P}}(1)$. Using $\sum_{i}a_{i}^{2}a_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}a_{i}^{2}\right)$ and Assumption S3 (f), we have
@@
so by Lemma S2, (S23) and Assumption S3 (a) and (b), (S24) holds.
Finally, combining (S20), (S24) with Lemmas S3 and S2 using Assumption S3 (a) and (b), we have that
@@
The conclusion now follows from Assumption S3 (a) and (b), Lemma S4 (b) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
S4.2.4 Auxiliary results
@@
S4.3 Robustness of nuisance function estimation for different estimators
S4.3.1 Estimation of nuisance functions with continuous $L$
S4.3.2 Estimation of nuisance functions with binary $L$
$Y:=2LW^{1}+\varepsilon.$
S4.3.3 Partially linear model estimate when partial linearity is violated
@@
are solutions to the minimization problem
@@
we obtain the desired result by minimizing the quadratic in $\theta$. ∎
S4.4 Computing directional perturbation effects in the presence of zero speeds
2:For $i\in[n]$, define $D_{i}:=\mathds{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
3:Define :=$\mathcal{I}:=\{i\in[n]\mid D_{i}=1\}$.
4:For $i\in\mathcal{I}$, define $(L_{i},W_{i})=\phi(Z_{i})$.
5:Compute estimate $\widehat{\tau}$ and asymptotic variance estimate $\widehat{\sigma}^{2}$ from either Algorithm S1 or S3.
6:Compute $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}$.
Theoretical guarantees for the proposed modification follow from repeating the arguments of the proof of Theorem S1 (a) and (c) to obtain the population quantities that determine the asymptotic behavior of $\widehat{\tau}$ and $\widehat{\sigma}^{2}$ and then applying Proposition S5 below.